A certain first-order reaction has a rate constant of 2.65×10−2 s−1 at 20 ∘C. What is the value of k at 50 ∘C if Ea = 70.0 kJ/mol ?
Answer #1

Acco Ling Ahhheniu 1ntes haki ny eura.hon log.kt. :3 Eszt [T.T ] .0 ohehe = Zo t 273-15 = 293, IS k = S0+ 273. IS =る23-15K 23